OpenStudy (anonymous):

Change the following mixed number to an improper fraction. 6 8/9

OpenStudy (anonymous):

and 41/6

OpenStudy (agreene):

\[6\frac{8}{9}=\frac{62}{9}\]

OpenStudy (anonymous):

62/9

OpenStudy (anonymous):

how about the 41/6?

OpenStudy (agreene):

\[4\frac{1}{6}=\frac{25}{6}\]

OpenStudy (anonymous):

the fraction is 41/6 not 4 1/6.

OpenStudy (agreene):

\[\frac{41}{6}=6\frac{5}{6}\]

OpenStudy (anonymous):

\[6+\left(\frac{8}{9}\right)=\frac{62}{9} \]agreene is correct.